OpenStudy (anonymous):

The perimeter of a rectangle is 34yd and the area of the rectangle is 60yd^2 Find the dimensions of the rectangle

OpenStudy (anonymous):

what are the answers?

OpenStudy (anonymous):

i need to know the answers can you help me

OpenStudy (anonymous):

is there any answers with the question?

OpenStudy (anonymous):

not that I know of.

OpenStudy (anonymous):

is it a 3D rectangle?

OpenStudy (anonymous):

No it doesn't say

OpenStudy (anonymous):

that's all the questions states

OpenStudy (anonymous):

 perimeter = 2x + 2y = 34 area = xy = 60 use substitution: from area we have y = 60/x so plug that in for y in the perimeter equation: 2x + 2(60/x) = 34 2x + 120/z = 34 multiply all by x 2x^2 -34x + 120 = 0 x^2 - 17x + 60 = 0 (x - 12)(x - 5) = 0 x is either 12 or 5 plugging those into xy = 60 y = either 12 or 5. that's your answer. either (x, y) = (12, 5) or (x. y) = (5, 12). both true

OpenStudy (anonymous):

pretty much one answer though: a 5 by 12 rectangle

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

glad i could help lol